OpenStudy (anonymous):

Kim walked to the grocery store and back home.The grocery store is 2 5/8 miles from her house. The next day kim walked twice as far to the post office and then back home.How far did she walk in all?

OpenStudy (texaschic101):

2 5/8 x 2 = 21/8 x 2 = 42/8 = 5 1/4 miles (first day) the 2nd day she walked twice as far.... 5 1/4 x 2 = 21/4 x 2 = 42/4 = 10 1/2 (2nd day) total : 5 1/4 + 10 1/2 = 21/4 + 21/2 = 21/4 + 42/4 = 63/4 = 15 3/4 miles ANSWER : Kim walked a total of 15 3/4 miles.